OpenStudy (paytorhater):

Find the missing term- (Blank) + (7 − 3i) + (5 + 9i) + 13i = 10 − 5i.

OpenStudy (soprano.h.d0816):

First, not to sound rude, but this would be a question for the Mathematics section of OpenStudy. If you have any questions, feel free to PM me or tag me in your post(:

OpenStudy (vamiter):

Well first you need to destribute in your parenthese 7-3i+5+9i+13i=10-5i then combine like terms 7+16i=10-5i Next plus 5i on both sides and minusing 7 on both sides 7-7+16i+5i=10-7-5+5=21i=3 now divide 0.14

OpenStudy (soprano.h.d0816):

Hey there!!! Since you are new here, read this tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 You might not have known, but you are in the "OpenStudy Feedback" section. Place your mouse over where it says OpenStudy Feedback, and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied! Please enjoy being on OpenStudy. And don’t forget to read OpenStudys’ Code of Conduct. This section is for providing your Feedback/Appreciation regarding the site and by Feedback, we mean "helpful information or criticism that is given to someone to say what can be done to improve a product( which is the site in this case)" You can use this section: ⇝ If you want to appreciate to OpenStudy ⇝ To share your thoughts on how to improve OpenStudy ⇝ To inform the site about any difficulty you are facing while using OpenStudy ⇝ To ask any other question related to this site. ⇝ If you want to post any posts such as congratulating someone, saying Happy Birthday, etc.